Suppose the currency-to-deposit ratio is 0.25, the excess reserve-to-deposit ratio is 0.03, and the required reserve ratio is 0.1. Which will have a larger impact on the money multiplier: a rise of 0.05 in the currency ratio or in the excess reserve ratio?

Instructions: Enter your response rounded to two decimal places.

If the currency-to-deposit ratio rises to 0.3, the multiplier will be m =

If, instead, the excess reserve-to-deposit ratio rises, the multiplier will be m =
Answer #1

Money multiplier, MM = (1 + Currency-deposit ratio) / (Currency-deposit ratio + Excess reserve ratio + Required Reserve ratio)

(a) Initially,

MM = (1 + 0.25) / (0.25 + 0.03 + 0.1) = 1.25 / 0.38 = 3.289

(b) Currency-deposit ratio = 0.3

MM = (1 + 0.3) / (0.3 + 0.03 + 0.1) = 1.3 / 0.45 = 2.89

(c) Excess reserve ratio rises to which number? MM cannot be computed unless exact number is provided.